OpenStudy (anonymous):

Find the arc length of y=2sqrt(x), with its starting point at (0,0)

terenzreignz (terenzreignz):

No arclength... unless you know where it ends... it has to end somewhere, right?

OpenStudy (anonymous):

It's just an integration from 0 to x. I'm stuck at the part of getting the integration done. \[\int\limits_{0}^{x} \sqrt(1+1/x) dx \]

terenzreignz (terenzreignz):

Maybe we shouldn't use x in the integrand... \[\Large \int\limits_0^x\sqrt{1+\frac1t}dt\]

terenzreignz (terenzreignz):

Anyway, currently, I'm stumped too... try integration by parts?

OpenStudy (turingtest):

I'm thinking trig sub\[x=\cot^2\theta\]

terenzreignz (terenzreignz):

Well, let's see where that goes :>

OpenStudy (anonymous):

i get \[t+1-1/2t^2lnt+1/4t^2\] from 0 to x with integration by parts, but I think it's nowhere close to the answer I have, which is \[\sqrt{x+x^2}+\ln(\sqrt(x)+\sqrt(1+x))\]

OpenStudy (anonymous):

oh nvm ive omitted the square root during my process, so i guess i'm out of luck :(

OpenStudy (turingtest):

trig sub worked out for me I think...\[x=\cot^2\theta\implies dx=-2\cot\theta\csc^2\theta d\theta=-2{\cos\theta\over\sin^3\theta}d\theta\]

OpenStudy (turingtest):

\[1+\cot^2\theta=\csc^2\]so our integral becomes... let me know if you see it from here

OpenStudy (turingtest):

\[\int\csc\theta\left(-2{\cos\theta\over\sin^3\theta}\right)d\theta=\cdots\]

OpenStudy (turingtest):

oh I messed up, it should have been tan in the radical

OpenStudy (anonymous):

yup i felt omething was weird too :)

OpenStudy (turingtest):

\[x=\cot^2\theta\implies dx=-2\cot\theta\csc^2\theta d\theta=-2{\cos\theta\over\sin^3\theta}d\theta\]\[\int\sqrt{1+\tan^2\theta}\left(-2{\cos\theta\over\sin^2\theta}\right)d\theta=\int\sec\theta\left(-2{\cos\theta\over\sin^2\theta}\right)d\theta\]

OpenStudy (turingtest):

I think that's better...

OpenStudy (turingtest):

\[=-2\int\csc^2\theta d\theta\]even easier

OpenStudy (anonymous):

so the result of the integration would be then -2cot(theta), but since we have x=cot^2 theta, how should we convert this back in terms of x?

OpenStudy (turingtest):

gotta draw the triangle

OpenStudy (anonymous):

i get sqrt(x)?

OpenStudy (turingtest):

I dunno, I'm just taking it as it comes... I didn't work out the final answer or anything

OpenStudy (turingtest):

I have feeling there was a way better way to do this :P

OpenStudy (anonymous):

hmm.. i guess we have usd other method or something since the derived solution looks far off from the actual solution I have, but thanks a bunch for helping me out :) i guess Ima leave this question for now :P

OpenStudy (turingtest):

dang, sorry, it's been a while since I've done this. may I ask what the actual solution is?

OpenStudy (turingtest):

oh I see you wrote it above

OpenStudy (anonymous):

its all good :) i will tackle this later on i guess :)

OpenStudy (turingtest):

how about\[u=1+\frac1x\implies du=-\frac1{x^2}dx\]\[dx=-x^2du=-\frac1{(u-1)^2}du\]

OpenStudy (zarkon):

I would use \[u=\sqrt{1+\frac{1}{t}}\] but that's just me :)

OpenStudy (turingtest):

Well @grit0614 that's you're answer, trust me. I'll try it out on my own too, good luck!

OpenStudy (zarkon):

using my substitution you will get to a point where you will need to use partial fractions.

OpenStudy (anonymous):

\[\int\limits_{0}^{x} \ -sqrt(u)*(1/(u-1)^2) du \] Wow this is getting nasty :) appreciate it. I will give it a try myself as well :)

OpenStudy (zarkon):

1) try my substitution 2) remember that your limits of integration change with u-sub

OpenStudy (anonymous):

I will try partial fraction as well :) thanks for all your help